# Application Management in C-ITS: Orchestrating Demand-Driven Deployments and Reconfigurations

### v1

@@ I Introduction
In future cooperative intelligent transport systems (C-ITS), various entities, such as vehicles equipped with driving automation systems, sensor-equipped roadside infrastructure units, edge/cloud servers, and control centers, will be connected, exchange data, and may offer computational resources. These advancements enable new applications – such as collective environment perception, cooperative decision-making, computation offloading, and intelligent traffic management – that can contribute to improved comfort and safety for road users [1], [2], [3]. Not only cloud and edge servers but also vehicles and roadside units can be part of a distributed computing system. However, these applications may also introduce complexity that is difficult to manage. The dynamic nature of C-ITS, the presence of resource-constrained entities, and the strict requirements for safety and security pose unique challenges.
Cloud-native techniques provide a promising foundation for the development and operation of scalable applications in dynamic environments. Such techniques involve paradigms like containerization, microservice architectures, and container orchestration. They enable loosely coupled systems which are manageable and resilient [4]. Said techniques and paradigms have the potential to contribute to the advancement of C-ITS.
Kubernetes has evolved as the de facto standard for orchestrating containerized applications in distributed systems. It is open-source and widely adopted by software companies worldwide [5]. Nevertheless, Kubernetes lacks methods that are domain-specific, e.g., to C-ITS, considering that specific tasks like the deployment of required applications are only needed at certain times or may depend on the specific content of data exchanged in the C-ITS. We have developed the approach RobotKube [6] to extend the regular capabilities of Kubernetes. RobotKube comprises software components designed to automate the identification of requirements and the formulation of specific Kubernetes workloads. These components include the event detector and the application manager.
@@
  • Complementation of RobotKube [6] by providing the concrete methodology of the application manager.
@@ II Related Work and Fundamentals
This work extends and further details our general orchestration approach introduced in RobotKube [6]. A central element of this approach is the operator application, formed by the application manager and the event detector with its operator plugin. Cluster operations in Kubernetes including the deployment and configuration of applications can be automated through operator applications. Operator applications act on events, which are occurrences of patterns in data, in the cluster. Events are detected by event detector components which can be implemented in the scope of a framework for event detection in C-ITS presented in [7]. As outlined in [6], the application manager receives a task description from the event detector’s operator plugin which the application manager then translates into a Kubernetes workload definition. The application manager composes the requested applications from available microservices and deploys them to nodes in the Kubernetes cluster. It is also capable of managing existing applications it has launched. The actual methodology how the application manager works is not yet described in [6]. The detailed methodology is one of the contributions of this paper.
@@
Our application management approach relies on concepts like containerization, microservices, and container orchestration. Such concepts are defined by the Cloud Native Computing Foundation (CNCF) as cloud-native techniques [4]. According to [8], containerization “is the process of packaging applications including their code and dependencies into single lightweight executables called container images”. Software platforms exist to run the process of containerization and the deployment of containers [9]. In microservices architectures, applications are built as a set of loosely coupled, “individual, independent (micro)services, with each service focused on a specific functionality” [10]. Separating functionality into distinct services makes them easier to deploy, update, and scale independently.
@@
Orchestration in the context of cloud-native techniques “refers to managing and automating the lifecycle of containerized applications in dynamic environments” [8]. This is executed through a container orchestrator enabling, e.g., deployments, scaling, auto-healing, and monitoring. A popular open-source container orchestrator is Kubernetes which provides software to build and deploy reliable, scalable distributed systems [5]. Kubernetes works declaratively, meaning that instead of describing how to perform operations, it is specified what the desired state of the system should be. Kubernetes then takes care of the details required to achieve this state. If the actual state of the cluster deviates from the desired one, Kubernetes controllers take action to reconcile the difference. [5], [11]
@@
Operators in terms of Kubernetes’ operator pattern are software extensions to Kubernetes that let the developer extend the cluster’s behavior without modifying the code of Kubernetes [12]. A Kubernetes operator is a specific kind of a controller with domain-specific logic implemented inside [13]. It allows more sophisticated logic defining what should happen when a deviation between the desired and the actual state of the cluster is detected. Our concept of custom operators is based on the operator pattern. We implement the custom operators using the framework Kopf [13].
@@
The need for orchestrating and managing services arises with vehicles becoming increasingly defined by software and with automotive service-oriented architectures (SOA) being developed [14], [15], [16]. Research is conducted with regard to in-vehicle orchestration [17], [18] and also to orchestration on vehicle and infrastructure level where intelligence can be distributed in a cluster of computing units in vehicles, edge, cloud, and infrastructure [19], [20]. Our work contributes to the orchestration on C-ITS level enabling the management of complex applications spanning across multiple nodes and involving multiple time-shifted requests with a demand-driven approach. In this paper, the term node refers to the context of cloud computing [21] and describes a physical computer which may be employed in a C-ITS entity.
@@
Figure 1: Reconciliation chain: An event detector [7] detects events (patterns in data) according to user-defined analysis rules and sends a deployment request containing a description of the current demand from the C-ITS. The application manager interprets the request, adds further application-specific configuration, and deploys Kubernetes custom resources (CRs) encoding the current demand on the application via the Kubernetes API. The custom operators process the declared demand encoded in the CRs, consult their internal bookkeeping, and formulate a desired state of the application. By deploying, reconfiguring, or shutting down the application, action is taken to align the current state with the desired state.
@@ III-A Design Principles for Applications
Applications are provided in the form of one or more container images built through automated containerization processes [22]. A container image is application-agnostic and can be reused in different applications. The Kubernetes resources necessary to run one or more containers in the cluster are packaged in Helm charts [23]. Helm charts offer parameterization allowing the deployment of containers to be adapted to the specific needs of the application. Container images and Helm charts are provided in application registries facilitating the integration of applications with continuous development processes. After development, an application has to pass certain steps of testing, including simulation [24], before it is made available for deployment.
@@ III-B Reconciliation Chain
Our application management approach employs the reconciliation chain extending the Kubernetes reconciliation loop where controllers constantly observe the current state of the cluster and take action to make the observed state match the desired state [5]. Although controllers automate the reconciliation of current and desired states, the desired state of each Kubernetes resource must be defined individually, e.g., by the user. The reconciliation chain automates the declaration of desired states of individual Kubernetes resources and extends the reconciliation loop to the application level enabling the demand-driven deployment, reconfiguration, and scaling of complex multi-service C-ITS applications requested by multiple C-ITS entities. The reconciliation chain, which is illustrated in Fig. 1, involves several components either developed by us or based on existing concepts of Kubernetes. The event detector with operator plugin [7], the application manager, and custom operators are the key components. Demands from the C-ITS are translated through the reconciliation chain into the desired state of an application, with subsequent actions taken to align the observed state with the desired state.
@@ III-C Application Manager
The application manager, together with an event detector with operator plugin [7], forms an operator application [6]. The event detector detects events, which are patterns in data exchanged in the C-ITS, according to implemented analysis rules. As a reaction upon an event, the event detector sends a deployment request to the application manager. It contains a formulated demand which may include information about which application is demanded, basic application configurations, necessary communication channels, and which C-ITS entities request the application. The event detector sends this information without having any knowledge about previous states of (running) applications, including whether it is already running or what configuration it may already be running with. The decision whether the application’s microservices are to be deployed, reconfigured, or shut down is made in the consecutive steps of the reconciliation chain.
@@ III-D Custom Operators
The subsequent step of the reconciliation chain is executed by the custom operators which process the declared demands encoded in the CRs, consult their internal bookkeeping to take past states into account, and formulate a new desired state of the application. Custom operators are implemented based on Kubernetes’ operator pattern as described in Section II. In particular, an operator watches for deviations between the current state of a Kubernetes custom resource (CR) and its declared desired state. If a deviation occurs, some developer-defined logic, e.g., the deployment of an application, is executed by the operator [12]. The content, e.g., the configuration of the application to be deployed, that is contained in a CR of a specific kind, is defined through a custom resource definition (CRD) [25]. The developer specifically designs a CRD to contain distinct domain-specific information based on which the operator does its work.
@@ IV Experimental Setup
As listed in Table I, the experimental setup entails various C-ITS entities, such as connected vehicles (CVs), a sensor-equipped roadside infrastructure station unit (RISU), an edge server, and a cloud server. The CVs send information about their current state, ego data [26], including their poses. CV $V_{0}$ and RISU $S$ are both equipped with lidar sensors and, therefore, provide point clouds. The setup is visualized in Fig. 2. All entities are represented as nodes in a Kubernetes cluster. On the nodes, microservices run in Kubernetes pods.
@@
The Kubernetes cluster is set up using k3d [27] which allows multi-node clusters while running on a single machine. The software components are based on ROS 2. ROS 2 bags recorded in simulation are used to simulate live data from the CVs and the RISU. MQTT is used for data transmission across nodes. To allow reproducibility, the experimental setup is made publicly available  [^1] https://github.com/ika-rwth-aachen/robotkube/
@@ ACKNOWLEDGMENT
This research is accomplished within the research projects “autotech.agil” (FKZ 1IS22088A) and “6GEM” (FKZ 16KISK036K). We acknowledge the financial support by the Federal Ministry of Research, Technology and Space (BMFTR).

### v2

@@ I Introduction
In future cooperative intelligent transport systems (C-ITS), various entities, such as vehicles equipped with driving automation systems, sensor-equipped roadside infrastructure units, edge/cloud servers, and control centers, will be connected, exchange data, and may offer computational resources. These advancements enable new applications – such as collective environment perception, cooperative decision-making, computation offloading, and intelligent traffic management – that can contribute to improved comfort and safety for road users [wang_homogeneous_2024], [liu_behavioral_2023], [kim_impact_2015]. Not only cloud and edge servers but also vehicles and roadside units can be part of a distributed computing system. However, these applications may also introduce complexity that is difficult to manage. The dynamic nature of C-ITS, the presence of resource-constrained entities, and the strict requirements for safety and security pose unique challenges.
Cloud-native techniques provide a promising foundation for the development and operation of scalable applications in dynamic environments. Such techniques involve paradigms like containerization, microservice architectures, and container orchestration. They enable loosely coupled systems which are manageable and resilient [cloud_native_computing_foundation]. Said techniques and paradigms have the potential to contribute to the advancement of C-ITS.
Kubernetes has evolved as the de facto standard for orchestrating containerized applications in distributed systems. It is open-source and widely adopted by software companies worldwide [burnsKubernetesRunningDive2019]. Nevertheless, Kubernetes lacks methods that are domain-specific, e.g., to C-ITS, considering that specific tasks like the deployment of required applications are only needed at certain times or may depend on the specific content of data exchanged in the C-ITS. We have developed the approach RobotKube [robotkube] to extend the regular capabilities of Kubernetes. RobotKube comprises software components designed to automate the identification of requirements and the formulation of specific Kubernetes workloads. These components include the event detector and the application manager.
@@
  • Complementation of RobotKube [robotkube] by providing the concrete methodology of the application manager.
@@ II Related Work and Fundamentals
This work extends and further details our general orchestration approach introduced in RobotKube [robotkube]. A central element of this approach is the operator application, formed by the application manager and the event detector with its operator plugin. Cluster operations in Kubernetes including the deployment and configuration of applications can be automated through operator applications. Operator applications act on events, which are occurrences of patterns in data, in the cluster. Events are detected by event detector components which can be implemented in the scope of a framework for event detection in C-ITS presented in [reiherevent]. As outlined in [robotkube], the application manager receives a task description from the event detector’s operator plugin which the application manager then translates into a Kubernetes workload definition. The application manager composes the requested applications from available microservices and deploys them to nodes in the Kubernetes cluster. It is also capable of managing existing applications it has launched. The actual methodology how the application manager works is not yet described in [robotkube]. The detailed methodology is one of the contributions of this paper.
@@
Our application management approach relies on concepts like containerization, microservices, and container orchestration. Such concepts are defined by the Cloud Native Computing Foundation (CNCF) as cloud-native techniques [cloud_native_computing_foundation]. According to [cncf_container_orchestration], containerization “is the process of packaging applications including their code and dependencies into single lightweight executables called container images”. Software platforms exist to run the process of containerization and the deployment of containers [matthias_docker_2016]. In microservices architectures, applications are built as a set of loosely coupled, “individual, independent (micro)services, with each service focused on a specific functionality” [cncf_microservices_architecture]. Separating functionality into distinct services makes them easier to deploy, update, and scale independently.
@@
Orchestration in the context of cloud-native techniques “refers to managing and automating the lifecycle of containerized applications in dynamic environments” [cncf_container_orchestration]. This is executed through a container orchestrator enabling, e.g., deployments, scaling, auto-healing, and monitoring. A popular open-source container orchestrator is Kubernetes which provides software to build and deploy reliable, scalable distributed systems [burnsKubernetesRunningDive2019]. Kubernetes works declaratively, meaning that instead of describing how to perform operations, it is specified what the desired state of the system should be. Kubernetes then takes care of the details required to achieve this state. If the actual state of the cluster deviates from the desired one, Kubernetes controllers take action to reconcile the difference. [burnsKubernetesRunningDive2019], [cncf_kubernetes]
@@
Operators in terms of Kubernetes’ operator pattern are software extensions to Kubernetes that let the developer extend the cluster’s behavior without modifying the code of Kubernetes [kubernetes_operator]. A Kubernetes operator is a specific kind of a controller with domain-specific logic implemented inside [kopf_concepts]. It allows more sophisticated logic defining what should happen when a deviation between the desired and the actual state of the cluster is detected. Our concept of custom operators is based on the operator pattern. We implement the custom operators using the framework Kopf [kopf_concepts].
@@
The need for orchestrating and managing services arises with vehicles becoming increasingly defined by software and with automotive service-oriented architectures (SOA) being developed [kampmann2022asoa], [kampmann_dynamic_2019], [rumezASOA2020]. Research is conducted with regard to in-vehicle orchestration [schindewolf_toward_2022], [nayak_automotive_2023] and also to orchestration on vehicle and infrastructure level where intelligence can be distributed in a cluster of computing units in vehicles, edge, cloud, and infrastructure [magaia_intelligent_2021], [arthurs_taxonomy_2022]. Our work contributes to the orchestration on C-ITS level enabling the management of complex applications spanning across multiple nodes and involving multiple time-shifted requests with a demand-driven approach. In this paper, the term node refers to the context of cloud computing [cncf_node] and describes a physical computer which may be employed in a C-ITS entity.
@@
Figure 1: Reconciliation chain: An event detector [reiherevent] detects events (patterns in data) according to user-defined analysis rules and sends a deployment request containing a description of the current demand from the C-ITS. The application manager interprets the request, adds further application-specific configuration, and deploys Kubernetes custom resources (CRs) encoding the current demand on the application via the Kubernetes API. The custom operators process the declared demand encoded in the CRs, consult their internal bookkeeping, and formulate a desired state of the application. By deploying, reconfiguring, or shutting down the application, action is taken to align the current state with the desired state.
@@ III-A Design Principles for Applications
Applications are provided in the form of one or more container images built through automated containerization processes [docker_ros]. A container image is application-agnostic and can be reused in different applications. The Kubernetes resources necessary to run one or more containers in the cluster are packaged in Helm charts [helm]. Helm charts offer parameterization allowing the deployment of containers to be adapted to the specific needs of the application. Container images and Helm charts are provided in application registries facilitating the integration of applications with continuous development processes. After development, an application has to pass certain steps of testing, including simulation [geller_carlos_2024], before it is made available for deployment.
@@ III-B Reconciliation Chain
Our application management approach employs the reconciliation chain extending the Kubernetes reconciliation loop where controllers constantly observe the current state of the cluster and take action to make the observed state match the desired state [burnsKubernetesRunningDive2019]. Although controllers automate the reconciliation of current and desired states, the desired state of each Kubernetes resource must be defined individually, e.g., by the user. The reconciliation chain automates the declaration of desired states of individual Kubernetes resources and extends the reconciliation loop to the application level enabling the demand-driven deployment, reconfiguration, and scaling of complex multi-service C-ITS applications requested by multiple C-ITS entities. The reconciliation chain, which is illustrated in Fig. 1, involves several components either developed by us or based on existing concepts of Kubernetes. The event detector with operator plugin [reiherevent], the application manager, and custom operators are the key components. Demands from the C-ITS are translated through the reconciliation chain into the desired state of an application, with subsequent actions taken to align the observed state with the desired state.
@@ III-C Application Manager
The application manager, together with an event detector with operator plugin [reiherevent], forms an operator application [robotkube]. The event detector detects events, which are patterns in data exchanged in the C-ITS, according to implemented analysis rules. As a reaction upon an event, the event detector sends a deployment request to the application manager. It contains a formulated demand which may include information about which application is demanded, basic application configurations, necessary communication channels, and which C-ITS entities request the application. The event detector sends this information without having any knowledge about previous states of (running) applications, including whether it is already running or what configuration it may already be running with. The decision whether the application’s microservices are to be deployed, reconfigured, or shut down is made in the consecutive steps of the reconciliation chain.
@@ III-D Custom Operators
The subsequent step of the reconciliation chain is executed by the custom operators which process the declared demands encoded in the CRs, consult their internal bookkeeping to take past states into account, and formulate a new desired state of the application. Custom operators are implemented based on Kubernetes’ operator pattern as described in Section II. In particular, an operator watches for deviations between the current state of a Kubernetes custom resource (CR) and its declared desired state. If a deviation occurs, some developer-defined logic, e.g., the deployment of an application, is executed by the operator [kubernetes_operator]. The content, e.g., the configuration of the application to be deployed, that is contained in a CR of a specific kind, is defined through a custom resource definition (CRD) [kubernetes_crd]. The developer specifically designs a CRD to contain distinct domain-specific information based on which the operator does its work.
@@ IV Experimental Setup
As listed in Table I, the experimental setup entails various C-ITS entities, such as connected vehicles (CVs), a sensor-equipped roadside infrastructure station unit (RISU), an edge server, and a cloud server. The CVs send information about their current state, ego data [perception_interfaces], including their poses. CV $V_{0}$ and RISU $S$ are both equipped with lidar sensors and, therefore, provide point clouds. The setup is visualized in Fig. 2. All entities are represented as nodes in a Kubernetes cluster. On the nodes, microservices run in Kubernetes pods.
@@
The Kubernetes cluster is set up using k3d [k3d] which allows multi-node clusters while running on a single machine. The software components are based on ROS 2. ROS 2 bags recorded in simulation are used to simulate live data from the CVs and the RISU. MQTT is used for data transmission across nodes. To allow reproducibility, the experimental setup is made publicly available  [^1] https://github.com/ika-rwth-aachen/robotkube/
@@ ACKNOWLEDGMENT
This research was conducted within the research projects “autotech.agil” (FKZ 1IS22088A) and “6GEM” (FKZ 16KISK036K), funded by BMFTR (Federal Ministry of Research, Technology and Space), and is continued in the project “iEXODDUS” (EU Horizon, GA No. 101146091).